OpenStudy (anonymous):

Simplify 7(2x+1)^3-(2x+1)^4

OpenStudy (anonymous):

your best bet is to write it as \[7(2x+1)^3-(2x+1)(2x+1)^3\] and then factor otherwise it will be a large mess

OpenStudy (anonymous):

7(2x^3+1)-(2x+1)(2x^3+1)

OpenStudy (anonymous):

right so far?

sam (.sam.):

Yeah you can also think of it like this, 7y^3-y^4 so y^3(7-y) (2x+1)^3(7-(2x+1)) (2x+1)^3(6-2x)

OpenStudy (anonymous):

7(2x^3+1)-(2x+1)(2x^3+1) 14x^3+7-(4x^4+2x+2x^3+1)

OpenStudy (anonymous):

im struggling with the right side

sam (.sam.):

From \(7(2x^3+1)-(2x+1)(2x^3+1)\) you missed the cubed for 2x+1, Should be like this, \[7(2x+1)^3-(2x+1)(2x+1)^3\] Simplifying or factoring, \[(2x+1)^3(7-(2x+1))\] \[(2x+1)^3(6-2x)\] Factor 2 out \[2(2x+1)^3(3-x)\]

OpenStudy (anonymous):

im not following u

OpenStudy (phi):

or, because I don't like (3-x), factor out -1 form (3-x) to get -1*(x-3) \[ -2(2x+1)^3(x-3) \]

OpenStudy (anonymous):

can u please show me step by step how to solve this problem

OpenStudy (phi):

7(2x^3+1)-(2x+1)(2x^3+1) right so far? no, this does not match the original you want 7(2x+1)^3 - (2x+1)(2x+1)^3 where you have the exponent of 3 on the everything in the parens

OpenStudy (anonymous):

what do u mean doesnt match the original?

OpenStudy (phi):

as posted above, if you rename (2x+1)^3 as y (so we can see the pattern) you would have 7*y -(2x+1)*y can you factor y from each term?

OpenStudy (anonymous):

ok i gotcha so u should look to have the same exponent

OpenStudy (anonymous):

wouldnt that be 7y-2x-1*y

OpenStudy (phi):

Let's start at the beginning 7(2x+1)^3-(2x+1)^4 rename (2x+1) as y (no exponent, so this is different from what I did up above) replace (2x+1) with y: 7*y^3 - y^4 follow?

OpenStudy (phi):

now write y^4 as y*y^3 (remember y^1 * y^3 = y^(1+3) = y^4 add exponents)

OpenStudy (anonymous):

yes i follow

OpenStudy (phi):

so rewrite y^4 as y*y^3 to get 6*y^3- y*y^3 y^3 is in both terms, so it can be "factored out" what do you get?

OpenStudy (anonymous):

u mean 7*y^3

OpenStudy (anonymous):

y^3(7-y)

OpenStudy (phi):

yes, 7y^3-y*y^3 looks good: y^3(7-y) now replace y with (2x+1) what do you get?

OpenStudy (anonymous):

(2x+1)^3(7-(2x+1))

OpenStudy (phi):

good. notice that (7-(2x+1)) is the same as (means) (7+ -1*(2x+1)) distribute the -1 by multiplying it times each term in the parens

OpenStudy (anonymous):

(2x+1)^3*(7-2x-1)

OpenStudy (phi):

now combine "like terms" in the 2nd parens: you have (reordering and making it all addition) +7 + -2x + -1 or -2x + 7 + -1

OpenStudy (anonymous):

(2x+1)^3*(6-2x)

OpenStudy (anonymous):

(2x^3+1)*(6-2x) 12x^3-4x^4+6-2x =-4x^4+12x^3-2x+6

OpenStudy (phi):

no, expanding this is messy. we can do that later (because you are not doing it correctly, and it is good to know how), but for now, the thing to do is factor -2 out of the last parens: (6-2x)

OpenStudy (anonymous):

(2x^3+1)*-2(6+x)

OpenStudy (phi):

lost the bubble, after all those correct steps! Here is where you were (2x+1)^3*(6-2x) first, do not do anything to the first term (2x+1)^3 for the other term, notice that both terms inside can be divided by -2: 6= -2*-3 and -2*x is -2*x: so (6+ -2x) is the same as (-2* -3 + -2*x) there is a -2 in each. can you factor it out?

OpenStudy (anonymous):

(2x+1)^3*(6-2x) (2x+1)^3*-2(-3+x)

OpenStudy (phi):

yes, now people put the -2 out front, rather than in the middle also, people like the x term first in that 2nd parens so, the standard way to write this is: -2(2x+1)^3 (x-3) follow?

OpenStudy (anonymous):

yes

OpenStudy (phi):

(-3+x) is the same as (x+ -3) (you can switch the order when adding) and (x+ -3) is generally written as (x-3) (looks simpler)

OpenStudy (anonymous):

okay

OpenStudy (phi):

-2(2x+1)^3 (x-3) is the answer though if it were multiple choice, you might see -2(x-3)(2x+1)^3 (different order, but the same thing, just like 2*3*4 is the same as 3*4*2)

OpenStudy (anonymous):

the answer should be 2(2x+1)^3(3-x)

OpenStudy (phi):

Now just a note on (2x+1)^3 the exponent 3, written as \[ (2x+1)^3 \] is short hand for multiplying the "base" (which is (2x+1) here) by itself 3 times. so, you could write \[ (2x+1)^3 =(2x+1)(2x+1)(2x+1)\]

OpenStudy (anonymous):

how would u go about multiplying that out

OpenStudy (phi):

you could use foil to multiply (2x+1)(2x+1) to get 4x^2+4x+1 then you have to multiply by (2x+1) again! ( 4x^2+4x+1)(2x+1) = 2x^3+8x^2+2x+4x^2+4x+1 which simplifies to 2x^3 + 12x^2 +6x+ 1 as you can see, it is messy. Now the trick is if you started with this 2x^3 + 12x^2 +6x+ 1 can you figure out it is (2x+1)^3 ?! (that is the hard to do!)